OpenStudy (anonymous):

Three fractions that are equivalent to: 12/36 4/5 and 3/12

OpenStudy (anonymous):

12/36 = 2/6 = 1/3 = 24/74 4/5 = 8/10 = 12/15 = 16/20 3/12 = 6/24 = 9/36 = 12/48